# DeepPrune: Parallel Scaling without Inter-trace Redundancy

### v1

@@ 5.4 Ablation Study
Besides, we analyze the trade-off between efficiency, answer diversity, and final accuracy by varying the redundancy threshold $\tau$ in DeepPrune (Table 3). The pass rate measures answer diversity after clustering, while accuracy reflects the voting outcome from the largest cluster. As $\tau$ decreases from 0.75 to 0.25, token consumption decreases significantly due to more aggressive pruning. However, this comes at the cost of reduced answer diversity, particularly on challenging problems. On AIME25, pass rate drops from 96.7% to 70% under greedy clustering, indicating that higher thresholds may prune valuable diverse reasoning paths. The majority voting accuracy shows the same trend. More analyses for this ablation are in Appendix E.
@@ 6 Conclusion
We identify inter-trace redundancy as a major efficiency bottleneck in parallel reasoning, where over 80% of computational resources are wasted on generating equivalent reasoning paths. To address this, we propose DeepPrune, a novel framework that trains a judge model for online greedy clustering to dynamically prune redundant traces while preserving answer diversity. Extensive experiments show that DeepPrune reduces token consumption up to 91.4% without hurting accuracy. Our work establishes that learned similarity judgment effectively addresses redundancy in parallel scaling, paving the way for more efficient reasoning systems.
@@ Appendix A Detailed Analysis of Inter-trace Redundancy
During the trace collection process, we initially sampled 16 responses for each query. However, some Chain-of-Thought traces failed to produce valid answers or did not terminate within the ‘max_length‘ limit. These invalid traces were excluded from the pair calculation. Furthermore, while our full dataset comprised total 758 problems from benchmarks including GPQA, AIME24, AIME25, and MATH500, due to computational resource constraints, we were only able to run all 758 queries on DeepSeek-R1-Distill-Llama-8B. For other models, we sampled 115 problems from the four datasets. This led to the following total pair counts and similarity ratios for each model in Table 4.
@@ Appendix C Hyperparameters
$\alpha_{t}$ 0.5 Focal loss class-balancing parameter
@@
$K$ 32 Maximum number of clusters
$K_{1}$ 10 Maximum number of sampled traces for cluster similarity calculation
$K_{2}$ 10 Max traces to finish reasoning from the largest cluster
$K_{3}$ 64 Number of traces to sample for reasoning if all clusters are singletons
@@ E.1 Comparison of Diversity with Baselines
Our approach, after clustering, theoretically retains distinct reasoning paths in different clusters. Each cluster’s representative trace could contribute to a pass@k calculation, where $k$ would typically be greater than one (equal to the number of clusters). This inherent diversity in our retained traces is evident in Table 3, where higher $\tau$ thresholds lead to higher pass rates under greedy clustering, indicating more diverse paths are preserved. To adaptively adjust the threshold, we tried setting the threshold according to the the cluster number. If the number of clusters exceed 16, then we increase the threshold by 0.03, until it reaches 0.9. Actually, this trick did not bring much changes to final accuracy and efficiency. So we did not conduct ablation on that. In conclusion, to ensure a fair comparison on a single-answer basis, our main online evaluation focuses on the final accuracy obtained after majority voting, as this is the metric that DeepConf and cons@512 also optimize for.

### v2

@@ 5.4 Ablation Study
Besides, we analyze the trade-off between efficiency, answer diversity, and final accuracy by varying the redundancy threshold $\tau$ in DeepPrune (Table 3). The pass rate measures answer diversity after clustering, while accuracy reflects the voting outcome from the largest cluster. As $\tau$ decreases from 0.75 to 0.25, token consumption decreases significantly due to more aggressive pruning. However, this comes at the cost of reduced answer diversity, particularly on challenging problems. On AIME25, pass rate drops from 96.7% to 70% under greedy clustering, indicating that higher thresholds may prune valuable diverse reasoning paths. The majority voting accuracy shows the same trend and the threshold $\tau=0.5$ provides the best balance.
@@ 6 Conclusion
We identify inter-trace redundancy as a major efficiency bottleneck in parallel reasoning, where over 80% of computational resources are wasted on generating equivalent reasoning paths. To address this, we propose DeepPrune, a novel framework that trains a judge model on out-of-distribution data for online greedy clustering to dynamically prune redundant traces while preserving answer diversity. Extensive experiments show that our OOD-trained judge model generalizes strongly to unseen reasoning models, achieving 0.7072 AUROC in offline evaluation, and DeepPrune reduces token consumption by up to 88.5% without hurting accuracy in online experiments. Our work establishes that learned similarity judgment effectively addresses redundancy in parallel scaling, paving the way for more efficient reasoning systems.
@@
Acknowledgements
This work is supported by the National Natural Science Foundation of China (625B2101, 62476150). This work is also supported by a grant from the Institute for Guo Qiang, Tsinghua University (2019GQB0003).
@@ Appendix A Detailed Analysis of Inter-trace Redundancy
During the trace collection process, we initially sampled 16 responses for each query. However, some Chain-of-Thought traces failed to produce valid answers or did not terminate within the ‘max_length‘ limit. These invalid traces were excluded from the pair calculation. Furthermore, while our full dataset comprised over 760 problems from benchmarks including GPQA, AIME24, AIME25, and Math500, due to computational resource constraints, we were only able to run all 760 queries on Deepseek-R1-Distill-llama-8b. For other models, we sampled approximately over 100 problems. This led to the following total pair counts and similarity ratios for each model in Table 4.
@@ Appendix C Hyperparameters
$\alpha_{t}$ 0.25 Focal loss class-balancing parameter
@@
$K$ 16 Maximum number of clusters
pp 20 Number of sampled traces for cluster similarity calculation
$q_{1}$ 20 Max traces to finish reasoning from the largest cluster
$q_{2}$ 32 Number of traces to sample for reasoning if all clusters are singletons
@@ E.1 Comparison of Diversity with Baselines
Our approach, after clustering, theoretically retains distinct reasoning paths in different clusters. Each cluster’s representative trace could contribute to a pass@k calculation, where $k$ would typically be greater than one (equal to the number of clusters). This inherent diversity in our retained traces is evident in Table 3, where higher $\tau$ thresholds lead to higher pass rates under greedy clustering, indicating more diverse paths are preserved. To ensure a fair comparison on a single-answer basis, our main online evaluation focuses on the final accuracy obtained after majority voting, as this is the metric that DeepConf and cons@512 also optimize for.